OpenStudy (anonymous):

show that the inflection points for y=sinx/x lie on the curve y^2(x^4 +4)=4

OpenStudy (phi):

find the second derivative wrt x of \[ y = \frac{\sin x }{x} \] and set y'' = 0 you will get an expression with sin x and cos x in it. from \( y = \frac{\sin x }{x}, \sin x = x y \) also, \( \cos x = \sqrt{1 - x^2 y^2}\) sub in those values. you will need to do a bit of algebra but you can (eventually) get y^2(x^4 +4)=4

OpenStudy (phi):

post how far you get.